OpenStudy (anonymous):

What is the vertex and the equation of the axis of symmetry of the graph of y=x^2-6x-7?

Nnesha (nnesha):

formula to find x -coordinate of the vertex \[\huge\rm x=\frac{ -b }{ 2a }\] then substitute all x for its value into the equation to find y-coordinate a=leading coefficient b= middle term

OpenStudy (anonymous):

so positive 6 over 2 times -1?

Nnesha (nnesha):

yes right !

Nnesha (nnesha):

wait -1 ?

Nnesha (nnesha):

it's 1x^2 so should be positive

OpenStudy (anonymous):

so (6/2(1))

Nnesha (nnesha):

right

Nnesha (nnesha):

so 6/2 =3 right ? now substitute x for 3 into the equation

OpenStudy (ryanwhite499):

Factor x2−6x−7 x2−6x−7 =(x+1)(x−7) Answer: (x+1)(x−7)